### v3

@@ 2.2 Network O&M Efficiency Improvement
For the continuous improvement and optimization of the performance of mobile communication networks, the introduction of AI has also played a crucial role in the relevant fields on the wireless communication network side, providing brand-new solutions to the problems in key areas such as data traffic prediction, BS energy-saving management, network parameter optimization, mobility management, and Resource Orchestration. AI can analyze a large amount of data and make real-time decisions, which is especially useful in wireless networks where many variables and parameters must be continuously monitored and adjusted to optimize system performance [64]. It can dynamically allocate resources such as bandwidth and power in wireless networks to optimize system performance.
In addition, the integration of AI enables the network to better understand and predict the complexity and dynamics of the network. Based on the prediction results, proactive and timely network maintenance can be carried out. Furthermore, by leveraging AI’s DL and pattern recognition technologies, a more in-depth understanding of network behavior can be achieved, realizing more optimal resource management and scheduling strategies. Subsequently, the role of AI in network O&M management will be introduced through specific cases.
@@ 2.2.1 AI-Based Traffic Prediction
The progressive deployment of 6G networks [65] heralds the increasing prevalence of new application scenarios like virtual reality (VR) and immersive communication. In such applications, the network traffic exhibits high variability and significant dynamics. In this context, precise and efficient network traffic prediction techniques have become particularly important. By predicting network traffic, it is possible to identify traffic fluctuations and peak periods in advance, thereby enabling dynamic allocation and scheduling of network resources.
In recent years, data-driven DL methods have gradually become the mainstream technology in the field of traffic prediction. These methods, leveraging the powerful modeling capabilities of DNNs, can effectively extract and capture potential feature relationships from complex and dynamically changing network environments. For instance, the authors of [66] employed recurrent neural network (RNN) to achieve joint spatiotemporal prediction, aiming to improve the accurate modeling of network traffic variations. The method proposed in [67], which is based on deep belief networks, effectively predicts the long-term variation trends of network traffic. The result in [68] demonstrated that, compared to conventional ̵autoregressive integrated moving average models, long short-yerm memory outperforms in traffic prediction tasks, particularly in handling nonlinear and complex time-series data.
In centralized traffic prediction methods, BS typically need to collect traffic data from various geographical locations, which inevitably introduces additional communication overhead and potential security risks. As a result, federated learning (FL) based traffic prediction methods have emerged. The FedDA method was proposed in [69] effectively reduces the delay and bandwidth overhead associated with data transmission by only transmitting model parameters instead of raw traffic data. This approach not only ensures privacy and security but also improves the efficiency of data processing. As shown in Figure 9, the authors of [70] proposed a distance-weighted federated traffic prediction method that better captures the spatiotemporal characteristics of traffic and improves the accuracy of traffic forecasting.
To reduce the model training time, the authors of [71] proposed a transfer learning-based approach, which accelerates the training process by uncovering the potential correlations between traffic patterns in different regions. Furthermore, other studies have employed techniques such as data sharing [72] and meta-learning [73] to further optimize the performance of traffic prediction models. In the future, network traffic prediction methods will focus not only on extracting spatiotemporal features and long-term dependencies but also on enhancing the lightweight design and privacy protection capabilities of the models.
Figure 9: Network traffic prediction
@@ 2.2.2 AI-Based BS Energy Conservation
According to statistics, the information and communication technology (ICT) industry directly contributes approximately 4% of global greenhouse gas emissions, with mobile communication networks accounting for over 10% of this figure. In response to the challenges posed by climate change, the ITU released Recommendation 1470, titled “Greenhouse Gas Emission Trajectories for the ICT Industry in Line with the United Nations Framework Convention on Climate Change Paris Agreement,” which calls for mobile operators to achieve a 45% reduction in carbon emissions from 2020 to 2030, aiming to drive the industry’s transition towards sustainability. Currently, the energy consumption of communication networks is primarily concentrated in computers and data centers, BS, and related network infrastructure, making energy-saving a critical issue that requires urgent attention.
@@
The energy optimization within 6G networks exhibits a close correlation with user latency. A key focus of academic research has been how to achieve network energy savings by effectively shutting down BS while ensuring the maintenance of basic service quality. The authors of [74] proposed a BS shutdown strategy based on traffic thresholds. However, during peak network periods, this approach may result in frequent state switching, leading to service interruptions. On the other hand, the authors of [75] classified BS states into four different sleep modes based on the activation time scale of the BS. While deeper sleep modes can significantly reduce energy consumption, they also result in increased wake-up delays, which may affect service quality.
@@
The energy-saving strategy for BS predicated on RL, via the incessant interactive learning between the agent and the operational milieu, dynamically modifies the parameters so as to accommodate the variations in service load and the oscillations in channel conditions. In [76], a BS state management scheme based on Q-learning was proposed. This approach dynamically adjusts the energy consumption of the BS in response to load changes by exploring the optimal duration of different sleep modes. The authors of [77] considered the interference issue between multiple BS and designs an energy optimization method based on collaborative strategies. By enabling cooperation between BS, this method reduces the overall system energy consumption. A cascading RL method was proposed in [78], which optimizes the decision process of user management and BS states in multi-BS scenarios, further improving network efficiency. The aforementioned studies primarily focus on internal strategy optimization within communication networks, with insufficient attention paid to the impact of external resources.
@@
Introducing renewable energy, especially clean energy sources such as solar power, as the energy supply for networks can significantly improve network energy efficiency and reduce dependence on conventional power grids. As depicted in Figure 10, the authors integrate RL with graph neural networks and accomplish the optimization of energy conservation and carbon emissions in wireless networks by capitalizing on photovoltaic solar technology [79]. Looking ahead, the energy-saving efforts in wireless networks should be deeply integrated with the use of renewable energy, providing more sustainable solutions.
Figure 10: Carbon emissions of different strategies
@@ 2.2.3 AI-Based Network Parameter Optimization
In wireless communication systems, parameter configuration has a profound impact on network performance. In particular, the rational setting of network parameters is crucial, as these parameters determine the design, energy consumption, and performance of the 6G network, and must be aligned with the real-time distribution of users and the electromagnetic environment. Based on the characteristics of wireless networks, network parameters can be categorized into two types: BS parameters and threshold parameters. Specifically, BS parameters involve basic switching functions such as VONR, paging, and slicing; threshold parameters configure the thresholds for cell reselection, event decision-making, and neighbor cell hand off. Moreover, spectrum resource planning in communication systems is also a critical component of parameter optimization. Studies have shown that spectrum is often underutilized in many time slots [80], and AI technologies have demonstrated great potential in cognitive radio [81]. AI methods can effectively monitor and mitigate interference [82], predict spectrum usage trends [83], and significantly improve the utilization of spectrum resources.
Conventional parameter optimization mainly relies on drive tests and expert experience. The current network parameter optimization methods are as follows: When network operators identify significant deficiencies in communication service quality, specific operators will deploy drive test vehicles to collect and report network performance metrics, such as network throughput and audio/video transmission delay. Subsequently, operators create scatter plots to further characterize network performance, and finally, expert experience is used to analyze and optimize the network parameters [84]. With the increasing complexity of network structures, conventional parameter optimization becomes inefficient. Due to the significant increase in the number of 6G antennas, the channel is no longer a simple one-dimensional circuit loss structure but has expanded into an N-dimensional angle power spectrum, resulting in a vast number of parameter combinations. The efficiency of expert-based parameter tuning is low, making it difficult to modify billions of 6G network parameters. Furthermore, conventional experience-based tuning methods often rely on inherent rules, which makes it difficult to fully capture and adapt to the complex relationships in the network [85]. Table 2 summarizes various parameters in the current networks.
@@
Data Type Data Name Application Area
Terminal data Block error rate Scheduling optimization
@@
Air interface data Multi-path time delay Channel modeling
@@
Core network data Operation log Fault handling
@@
Service data Online duration Personalized service
@@
Currently, AI models have demonstrated unprecedented advantages in large-scale parameter tuning, nonlinear fitting, and real-time optimization [86]. Specifically, the increase in the number of antennas and the complexity of the 6G network has led to a surge in network parameter data, and AI can quickly search and optimize these combinations. In the context of the highly intricate architecture and nonlinear traits of the 6G network [87], AI technology can gain a more profound understanding of the network’s nonlinear features by learning from and dissecting extensive network data, thereby determining a more optimal parameter configuration. Moreover, in the 6G network, user requirements and network conditions change swiftly. AI technology can promptly optimize network parameters based on real-time data and scenario necessities, attaining enhanced performance.
In recent related research arenas, deep RL (DRL) has been deployed for the optimization of network parameters [88]. Specifically, each performance metric within the current communication network is considered as a distinct RL state, and the adjustment operations carried out on network parameters are designated as diverse actions. In this procedure, by incessantly adapting the behavior of the RL model across different states, the model will receive feedback information from the environment to assess the merit of each action. RL models typically focus on reinforcing those actions that can yield higher reward returns. Once the reward cost stabilizes, DRL will progressively approach a convergent state and subsequently determine the optimal configuration scheme of wireless network parameters.
@@ 2.2.4 AI-Based Mobility Management
Mobility management is a fundamental function of mobile communication systems, encompassing the network’s management of mobile UE location information, security, and service continuity. Conventional mobility management typically functions by depending on a set of meticulously pre-configured policies and rigid rules. Nevertheless, in practical application contexts, these conventional mobility management strategies have progressively manifested a multitude of bottleneck issues and are incapable of efficiently addressing the intricate and variable 6G network environment as well as the diversified demand circumstances of users. Specifically, when the number of mobile users increases rapidly, these methods may not be able to quickly and effectively handle interactions between users, leading to excessive network load and decreased performance. Second, in complex environments such as large public spaces or major events, conventional mobility management approaches may struggle to cope, resulting in network congestion. Additionally, for high-speed users, such as those in vehicles, frequently hand off can lead to unstable network connections.
To effectively transcend these challenges, AI technology, leveraging its robust data processing capabilities, remarkable adaptive learning mechanisms and highly automated decision-making and optimization capabilities, is integrated into the domain of mobility management. AI is capable of conducting in-depth data mining and analysis of extensive network data, such as user behavior data, network traffic data and channel quality data, to extract valuable insights and latent regular patterns. Additionally, it can automatically and dynamically adjust and continuously optimize mobility management strategies without the need for manual human interference, relying on the acquired knowledge. As early as 2015, 3GPP began researching how AI technology could be integrated into the 5G network. For example, the 3GPP SA2 and SA5 working groups (WG) launched projects on “Enablers for Network Automation” and “Enhancement of Management Data Analytics Services”, which made good progress [89].
@@
Specifically, by means of technologies such as DL, AI is capable of predicting user behaviors including moving speed and direction, thereby permitting the proactive selection and hand off between cells to diminish signal overhead. For example, in ultra-dense network scenarios, frequent cell exchanges can lead to increased overhead and connection disruptions, adversely affecting QoS. The authors of [90] proposed the use of a GRU-based RNN to adaptively handle hand off decisions, significantly reducing hand off latency and improving network throughput. In [91], the authors suggested a method for predicting the next position of vehicles using AI in ultra-dense network scenarios, which allows for BS selection based on predicted positions, thereby reducing hand off frequency and failure rates, and saving hand off signaling overhead. This demonstrates that AI-based mobility management is feasible and practical.
@@
On the other hand, through RL and other techniques, AI can dynamically optimize mobility management strategies based on network environment and user behavior, better addressing complex network environments. As wireless networks at higher frequencies require more dense deployment of BS, the authors of [92] addressed the issue of reduced network capacity due to frequently hand offs, proposing a centralized and multi-agent deep Q-network (DQN) algorithm. Both algorithms can find near-optimal user-to-BS allocations. As shown in Figure 11, the proposed DQN algorithm significantly reduces or even completely avoids ping-pong hand offs compared to the baseline algorithm (conventional SINR-based hand off algorithm). The authors of [93] proposed a ML-based framework to optimize the hand off between sub-6 GHz and millimeter wave bands in vehicular networks, using RL to enable vehicles to predict and discover the optimal hand off target in real time. The proposed method reduces hand off latency, improves connectivity in vehicular networks, achieves seamless hand off even in high mobility scenarios, and enhances user experience.
@@
In light of the aforementioned analysis, We can observe that mobility management is an eternal and significant topic in wireless networks. Most conventional solutions were designed and validated in previous network generations and may not meet future demands for higher communication quality and user experience. AI-based mobility management, including advanced and efficient self-optimization of cell selection, load balancing, and hand off parameters, etc., will play a more critical role in future 6G networks.
@@ 2.2.5 AI-Based Resource Orchestration
In the context of the deep fusion between 6G communication networks and the IoT, with hundreds of billions of devices being connected, the generated data is undergoing a continuous explosive growth [94]. Merely strengthening the communication capabilities is insufficient to meet the real-time requirements for data processing in control application scenarios [95]. By harnessing the communication and computing capabilities at the network edge and on the device side, the entire process of data collection, processing, analysis, and decision-making can be accomplished closer to the device side, avoiding the latency drawbacks caused by congestion in the CN, and thus significantly enhancing the data processing response speed. In light of the high-dimensional characteristics and unpredictability of the system resource state space, AI-based resource orchestration strategies (including RL and DRL) can attune to the alterations in the system state via continuous interaction with the system environment, thereby augmenting the overall performance of the system.
@@
The joint scheduling of communication and computing resources founded on RL attains the collaborative and optimal allocation of communication and computing resources via the interaction and learning between the agent and the 6G network environment [96]. Confronted with the resource allocation difficulties within the space-air-ground integrated network in the 6G era, the authors of [97] put forward the RL-assisted bandwidth aware virtual network resource allocation algorithm. This algorithm uses RL with a policy network for node embedding, preferentially handling high-bandwidth requests to meet strict user bandwidth demands. Given the relevant problems that exist within 6G mobile devices and communication systems, motivated by the concept of virtual network embedding, the authors in [98] pioneered a multi-objective aware dynamic resource scheduling algorithm designed for multi-layer computing networks, aiming to augment resource flexibility. They established a scheduling network underpinned by DRL and refined the learning process, thus proffering a sustainable resolution for resource scheduling strategies. Regarding the circumstance that the applicability of the open radio access network (O-RAN) in governing and optimizing the functions of the wireless access network within the 6G network has not been extensively explored, the authors of [99] developed low-complexity algorithms based on methods such as RL to address the problems of jointly optimizing traffic splitting and allocation, congestion control, and scheduling across different time scales. This effort provides insights for realizing a fully automated network with enhanced control and flexibility.
@@
On the other hand, the dynamic scheduling of network slices based on DRL exploits the potent feature extraction and learning capabilities of DNNs to manage network slice resources intelligently. To address the challenges of computing resource demands and intelligent resource allocation in 6G network IoT services, the authors of [100] proposed an effective DRL-based solution for IoT resource expansion and service placement, and verified the effectiveness of its multi-application autonomous resource allocation via dataset simulation. According to [101], the authors proposed a multi-task DRL method based on graph convolutional networks by introducing a joint network slicing and routing mechanism, achieving robustness in various network environments. In the study of [102], the authors established a model-free DRL framework by a hierarchical structure integrating modified deep deterministic policy gradient (DDPG) and double DQN to actualize an intelligent RAN slicing strategy with two-layer control granularity for maximizing QoS and slice spectrum efficiency. Moreover, a dynamic spectrum allocation scheme was proposed in [103], which utilized the advantages of the DQN and reduced the search state explosion by a reward and punishment framework to dynamically allocate unallocated resource blocks to mobile units and obtained the Pareto optimal solution of sub-problems via Chebyshev decomposition.
@@
The AI-based resource orchestration has already demonstrated certain advantages in fields such as 6G and has achieved remarkable results in resource allocation and utilization by combining various techniques. However, it still faces numerous challenges and opportunities in the future. On the one hand, efforts should be focused on optimizing and innovating algorithms. For instance, the improvement of DRL algorithms is urgent in order to cope with complex and changeable environments. On the other hand, when deepening the integration of multiple technologies, not only should the combination paths with emerging technologies be actively explored, but also the compatibility issues during the integration process should be addressed. Meanwhile, it is necessary to strive to enhance the adaptability to complex scenarios, meet the diverse resource requirements in various special scenarios of different industries, and improve performance evaluation and enhance algorithm interpretability, thereby achieving more extensive applications and development.

### v4

@@ 2.2 Network O&M Efficiency Improvement
AI can analyze a large amount of data and make real-time decisions, which is especially useful in wireless networks where many variables and parameters must be continuously monitored and adjusted to optimize system performance [68]. It can dynamically allocate resources such as bandwidth and power in wireless networks to optimize system performance. In addition, the integration of AI enables the network to better understand and predict the complexity and dynamics of the network. Based on the prediction results, proactive and timely network maintenance can be carried out. By leveraging AI’s DL and pattern recognition technologies, a more in-depth understanding of network behavior can be achieved, realizing more optimal resource management and scheduling strategies. Subsequently, the role of AI in network O&M management will be introduced through specific cases.
@@ 2.2.1 AI-Based Traffic Prediction
The progressive deployment of 6G networks [69] heralds the increasing prevalence of new application scenarios like virtual reality (VR) and immersive communication. In such applications, the network traffic exhibits high variability. By predicting network traffic, it is possible to identify traffic fluctuations and peak periods in advance, thereby enabling dynamic allocation and scheduling of network resources. For instance, the authors of [70] employed a recurrent neural network (RNN) to achieve joint spatiotemporal prediction, aiming to improve the accurate modeling of network traffic variations. The method proposed in [71], based on deep belief networks, effectively predicts the long-term variation trends of network traffic. The result in [72] demonstrated that, compared to conventional ̵autoregressive integrated moving average models, long short-term memory outperforms in traffic prediction tasks, particularly in handling nonlinear and complex time-series data.
In centralized traffic prediction methods, a BS typically needs to collect traffic data from various geographical locations, which inevitably introduces additional communication overhead and potential security risks. As a result, federated learning (FL)-based traffic prediction methods have emerged. The FedDA method was proposed in [73], which effectively reduces the delay and bandwidth overhead associated with data transmission by only transmitting model parameters, instead of raw traffic data. This approach not only ensures privacy but also improves the efficiency of data processing. The authors of [74] proposed a distance-weighted federated traffic prediction method that better captures the spatiotemporal characteristics of traffic and improves the accuracy of traffic forecasting.
To reduce the model training time, the authors of [75] proposed a transfer learning-based approach, which accelerates the training process by uncovering the potential correlations between traffic patterns in different regions. Other studies have employed techniques such as data sharing [76] and meta-learning [77], to further optimize the performance of traffic prediction models. In the future, network traffic prediction methods will extract spatiotemporal features and long-term dependencies, but also enhance the lightweight design and privacy protection capabilities of the models.
@@ 2.2.2 AI-Based BS Energy Conservation
The information and communication technology (ICT) industry directly contributes approximately 4% of global greenhouse gas emissions, with mobile communication networks accounting for over 10% of this figure [78]. In response to the challenges posed by climate change, the ITU released Recommendation 1470, titled “Greenhouse Gas Emission Trajectories for the ICT Industry in Line with the United Nations Framework Convention on Climate Change Paris Agreement”, which calls for mobile operators to achieve a 45% reduction in carbon emissions from 2020 to 2030, aiming to drive the industry’s transition towards sustainability.
@@
The energy optimization within 6G networks exhibits a close correlation with user latency. A key focus of academic research has been how to achieve network energy savings by effectively shutting down BS while ensuring the maintenance of basic service quality. The authors of [79] proposed a BS shutdown strategy based on traffic thresholds. During peak network periods, this approach may result in frequent state handoffs, leading to service interruptions. The authors of [80] classified BS states into four different sleep modes based on the activation time scale of the BS. While deeper sleep modes can significantly reduce energy consumption, they also result in increased wake-up delays, affecting service quality.
Figure 9: Carbon emissions of different strategies [81]
@@
The energy-saving strategy for BS predicated on reinforcement learning (RL), via the incessant interactive learning between the agent and the operational milieu, dynamically modifies the parameters to accommodate the variations in service load and the oscillations in channel conditions. In [82], a BS state management scheme was proposed based on Q-learning. It dynamically adjusts the energy consumption of the BS in response to load changes by exploring the optimal duration of different sleep modes. The authors of [83] considered the interference between multiple BS, and designed an energy optimization method based on collaborative strategies. By enabling cooperation between BS, this method reduces the overall system energy consumption. A cascading RL method was proposed in [84], which optimizes the decision process of user management and BS states in multi-BS scenarios, further improving network efficiency.
@@
Introducing renewable energy, especially clean energy sources, e.g., solar power, as the energy supply for networks can significantly improve network energy efficiency and reduce dependence on conventional power grids. As depicted in Figure 9, the authors of [81] integrate RL with graph neural networks and accomplish the optimization of energy conservation and carbon emissions in wireless networks by capitalizing on photovoltaic solar technology. Energy-saving efforts in wireless networks should be deeply integrated with the use of renewable energy, providing more sustainable solutions.
@@ 2.2.3 AI-Based Network Parameter Optimization
The rational setting of network parameters is crucial, as these parameters determine the design, energy consumption, and performance of the 6G network, and must be aligned with the real-time distribution of users and the electromagnetic environment. Wireless network parameters can be categorized into two types: BS parameters and threshold parameters. The BS parameters involve basic handoff functions such as paging and slicing. The threshold parameters configure the thresholds for cell reselection, event decision-making, and neighbor cell handoff. Moreover, spectrum resource planning in communication systems is a critical component of parameter optimization. Studies have shown that spectrum is often underutilized [85], and AI technologies have demonstrated great potential in cognitive radio [86]. AI can monitor and mitigate interference [87], predict spectrum usage trends [88], and significantly improve the utilization of spectrum resources.
Conventional parameter optimization relies mainly on drive tests and expert experience. When network operators identify significant deficiencies in communication service quality, specific operators will deploy drive test vehicles to collect and report network performance metrics, such as network throughput and audio/video transmission delay. Subsequently, operators create scatter plots to further characterize network performance. Finally, expert experience is used to analyze and optimize the network parameters [89]. Due to the significant increase in the number of 6G antennas, the channel is not a simple one-dimensional circuit loss structure but has expanded into an $N$-dimensional angle power spectrum. The efficiency of expert-based parameter tuning is low, making it difficult to modify billions of 6G network parameters. Furthermore, conventional experience-based tuning methods often rely on inherent rules, which makes it difficult to fully capture and adapt to the complex relationships in the network [90]. Table 2 summarizes various parameters in the current networks.
@@
Parameter Type Parameter Name Application Area
Terminal parameter Block error rate Scheduling optimization
@@
Air interface parameter Multi-path time delay Channel modeling
@@
Core network parameter Operation log Fault handling
@@
Service parameter Online duration Personalized service
@@
AI models have demonstrated unprecedented advantages in large-scale parameter tuning, nonlinear fitting, and real-time optimization [91]. In the context of the highly intricate architecture and nonlinear traits of the 6G networks [92], AI technology can gain a more profound understanding of the network’s nonlinear features by learning from and dissecting extensive network data, and optimizing network parameters based on real-time data and scenario. Deep RL (DRL) has been deployed for the optimization of network parameters [93]. Specifically, each performance metric within the current communication network is considered a distinct RL state, and the adjustment operations carried out on network parameters are designated as diverse actions. In this procedure, by incessantly adapting the behavior of the RL model across different states, the model will receive feedback information from the environment to assess the merit of each action. RL models typically reinforce those actions that can yield higher reward returns.
@@ 2.2.4 AI-Based Mobility Management
As early as 2015, 3GPP began researching how AI technology could be integrated into the 5G network. For example, the 3GPP SA2 and SA5 working groups (WG) launched projects on “Enablers for Network Automation” and “Enhancement of Management Data Analytics Services”, which made good progress [94]. Using technologies such as DL, AI is capable of predicting user behaviors, including moving speed and direction, thereby permitting the proactive selection and handoff between cells to diminish signal overhead. In ultra-dense network scenarios, frequent cell exchanges can lead to increased overhead and connection disruptions, adversely affecting QoS. The authors of [95] proposed the use of a GRU-based RNN to adaptively handle handoff decisions, significantly reducing handoff latency and improving network throughput. In [96], the authors suggested a method for predicting the next position of vehicles using AI in ultra-dense network scenarios, which allows for BS selection based on predicted positions, thereby reducing handoff frequency and failure rates, and saving handoff signaling overhead. This demonstrates that AI-based mobility management is feasible and practical.
@@
On the other hand, through RL and other techniques, AI can dynamically optimize mobility management strategies based on network environment and user behavior, better addressing complex network environments. As wireless networks in higher frequencies require more dense deployment of BS, the authors of [97] addressed the issue of reduced network capacity due to frequent handoffs, proposing a centralized and multi-agent deep Q-network (DQN) algorithm. Both algorithms can find near-optimal user-to-BS allocations. As shown in Figure 10, the DQN algorithm significantly reduces or even completely avoids ping-pong handoffs compared to the baseline (conventional signal-to-interference-plus-noise ratio (SINR)-based handoff algorithm). The authors of [98] proposed an ML-based framework to optimize the handoff between sub-6 GHz and mmWave bands in vehicular networks, using RL to enable vehicles to predict and discover the optimal handoff target in real-time. The method reduces handoff latency, improves connectivity in vehicular networks, achieves seamless handoff even in high mobility scenarios, and enhances user experience.
@@ 2.2.5 AI-Based Resource Orchestration
With hundreds of billions of devices being connected, the generated data is undergoing continuous explosive growth [99]. Merely strengthening the communication capabilities is insufficient to meet the real-time requirements for data processing in control application scenarios [100]. By harnessing the communication and computing capabilities at the network edge and on the device side, the entire process of data collection, processing, analysis, and decision-making can be accomplished closer to the device side, avoiding the latency drawbacks caused by congestion in the CN.
@@
The joint scheduling of communication and computing resources founded on RL attains the collaborative and optimal allocation of communication and computing resources via the interaction and learning between the agent and the 6G network environment [101]. Confronted with the resource allocation difficulties within the space-air-ground integrated network in the 6G era, the authors of [102] put forward an RL-assisted bandwidth-aware virtual network resource allocation algorithm. This algorithm uses RL with a policy network for node embedding, preferentially handling high-bandwidth requests to meet strict user bandwidth demands. Given the relevant problems that exist within 6G mobile devices and communication systems, motivated by the concept of virtual network embedding, the authors of [103] pioneered a multi-objective aware dynamic resource scheduling algorithm designed for multi-layer computing networks, aiming to augment resource flexibility. They established a scheduling network underpinned by DRL and refined the learning process, thus proffering a sustainable resolution for resource scheduling strategies. Regarding the circumstance that the applicability of the open radio access networks (O-RAN) in governing and optimizing the functions of the wireless access network within the 6G network has not been extensively explored, the authors of [104] developed low-complexity algorithms based on methods such as RL to address the problems of jointly optimizing traffic splitting and allocation, congestion control, and scheduling across different time scales. This effort provides insights for realizing a fully automated network with enhanced control and flexibility.
@@
To address the challenges of computing resource demands and intelligent resource allocation in 6G network IoT services, the authors of [105] proposed an effective DRL-based solution for IoT resource expansion and service placement, and verified the effectiveness of its multi-application autonomous resource allocation via dataset simulation. In [106], the authors proposed a multi-task DRL method based on graph convolutional networks by introducing a joint network slicing and routing mechanism, achieving robustness in various network environments. In the study of [107], the authors established a model-free DRL framework by a hierarchical structure integrating modified deep deterministic policy gradient (DDPG) and double DQN to actualize an intelligent RAN slicing strategy with two-layer control granularity for maximizing QoS and slice spectrum efficiency. Moreover, a dynamic spectrum allocation scheme was proposed in [108], which utilized the advantages of the DQN and reduced the search state explosion by a reward and punishment framework to dynamically allocate unallocated resource blocks to mobile units and obtained the Pareto optimal solution of sub-problems via Chebyshev decomposition.
@@
The AI-based resource orchestration still faces numerous challenges and opportunities in the future. For instance, the improvement of DRL algorithms is urgent to cope with complex and changeable environments. It is necessary to strive to enhance the adaptability to complex scenarios, meet the diverse resource requirements in various special scenarios of different industries, and improve performance evaluation and enhance algorithm interpretability.